OpenStudy (anonymous):

Figure out the total price. Original cost: $3.45 Markup: 29% Tax: 2%

OpenStudy (anonymous):

Please explain too, I need to figure out how to do the rest of these too

OpenStudy (anonymous):

Does tax go on the markup cost too, or just the original cost?

OpenStudy (anonymous):

I don't know, it just says Find the selling price of each item. :P

OpenStudy (anonymous):

Ok, I'm going to assume that tax goes on the markup too. Let's add markup first: 29% of 3.45 is 0.29*3.45=1.00

OpenStudy (anonymous):

So add that to the original price 3.45+1.00=4.45 Now for the tax: 0.02*4.45=0.09 add that to what we had: 4.45+0.09=4.54 So 4.54 is the final cost.

OpenStudy (anonymous):

My friend (in the same class) says you have to do cost times markup plus tax times cost. I dont know, Im super confused.

OpenStudy (anonymous):

If that helped any :P

OpenStudy (anonymous):

Ok I'll go with that then :T thanks!